OpenStudy (anonymous):

The area of a square is 4x^4/9y^2 sq units. What is the length of one side, in units. A. 18x^2/18y B. 4x/9y C. 2x^2\3y D. 4x^2/ 6y^2

mathslover (mathslover):

Area of a square is given by : \(\textbf{(side)}^2\) It is given to you that : Area of Square = \((\text{side})^2 = \dfrac{4x^4}{9y^2} sq. units \) To find the side of the square : Square root both sides : \(\sqrt{(\text{side})^2} = \sqrt{\dfrac{4x^4}{9y^2} sq. units }\) Thus, side = \(\sqrt{\dfrac{(2x^2)^2}{(3y)^2} } units\) Can you simplify it further?

OpenStudy (anonymous):

Yes , Thank you. :)

mathslover (mathslover):

You're welcome :)